# What’s under the hood: Investigating Automatic Metrics on Meeting Summarization

### v1

@@ 7 Conclusion
Our study, while offering a comprehensive analysis of challenges and errors in meeting summarization, primarily focuses on English-language summaries. This linguistic focus may lead to variations in challenges and errors across languages. Some challenges, especially those subtle to human perception, might be overlooked, creating potential gaps in our annotations. Our reliance on Google Scholar, despite its broad coverage, has its criticisms, as it tends to favor citation counts Fagan (2017) and may include less reputable sources Beall (2016). The QMSum dataset, with its 35 samples, provides statistical significance but represents only a fraction of potential meeting types. Thus, findings may vary across different meeting contexts. The challenges we associate with QMSum are inferred from human annotations rather than directly tested, and our model selection, while reflective of the current landscape, doesn’t capture every variant. Our choice of encoder-decoder models is comprehensive, but we miss out on architectures like hierarchical models due to accessibility issues which hamper the comparability of the models. The array of autoregressive models is continually expanding, and our selection is based on the current standings on the Huggingface LLM Leaderboard. Our method of linking challenges to errors is holistic, yet it might dilute strong connections between them, as we did not test the linking with isolated challenges. Certain metrics, like LENS and QuestEval, may have biased scores since they use the meeting transcript as input and aren’t domain-trained. Lastly, our findings, rooted in the nuances of meeting summarization, might not seamlessly apply to broader summarization domains like dialogue summarization, given each domain’s distinctive traits.
@@
Our research abides by ethical guidelines for AI research and is committed to privacy, confidentiality, and intellectual property rights. We’ve ensured that the datasets in our study, publicly available, do not house sensitive or personal details. While our study leverages existing resources and generative models, it’s important to note that these models can possess biases and may occasionally generate summaries with distortions, biases, or inappropriate content Gooding (2022). To counteract this, we’ve configured our models to omit potentially harmful or unsafe content. While our research aims to enhance meeting summarization to benefit communication and productivity across sectors, we’re acutely aware of the ethical challenges posed by AI in this domain. Meeting summarization models must be wielded with respect to privacy and consent, especially when processing sensitive or confidential material. It’s paramount that these models neither violate privacy nor perpetuate harmful biases. As the field evolves, we stress the importance of maintaining these ethical considerations and encourage fellow researchers to uphold them, ensuring that AI advancements in meeting summarization are both beneficial and ethically grounded. An integral aspect of our ethical commitment is reflected in our approach to annotator recruitment and management. The team of annotators, consisting of interns, student assistants, and doctoral students, was meticulously selected through internal channels. This strategy was chosen to uphold a high standard of annotation quality—a quality we found challenging to guarantee through external platforms such as Amazon Mechanical Turk. Ensuring fair compensation, these annotators were remunerated in accordance with institutional guidelines for their respective positions. Further, flexibility in the annotation process was also a priority. Annotators had the freedom to choose their working times and environments to prevent fatigue from affecting their judgment.
8 Bibliographical References

### v2

@@ Acknowledgements
Our study, while offering a comprehensive analysis of challenges and errors in meeting summarization, primarily focuses on English-language summaries. This linguistic focus may lead to variations in challenges and errors across languages. Some challenges, especially those subtle to human perception, might be overlooked, creating potential gaps in our annotations. Our reliance on Google Scholar, despite its broad coverage, has its criticisms, as it tends to favor citation counts Fagan (2017) and may include less reputable sources Beall (2016).
The QMSum dataset, with its 35 samples, provides statistical significance but represents only a fraction of potential meeting types. Thus, findings may vary across different meeting contexts.
We include a total of 175 samples, comparable to the original QMSum dataset (i.e., 232 samples), one of the most used datasets in meeting summarization. The considered dataset maintains a diverse representation of meeting types (academic, business, parliament) within the dataset. We acknowledge the limitation that QMSum might not cover all meeting types in existence. However, we focus on creating a unified understanding of the challenges in meeting summarization and the errors that occur when these challenges are unmet. For that, we used QMSum as a proxy.
The challenges we associate with QMSum are inferred from human annotations rather than directly tested, and our model selection, while reflective of the current landscape, does not capture every variant. Our choice of encoder-decoder models is comprehensive, but we miss out on architectures like hierarchical models due to accessibility issues, which hamper the comparability of the models. The array of large language models is continually expanding, and our selection is based on the standings on the Huggingface LLM Leaderboard at the time of writing. Our method of linking challenges to errors is holistic, yet it might dilute strong connections between them, as we did not test the linking with isolated challenges. Specific metrics, like LENS and QuestEval, may have biased scores since they use the meeting transcript as input and are not domain-trained. Lastly, our findings, rooted in the nuances of meeting summarization, might not seamlessly apply to broader summarization domains like dialogue summarization, given each domain’s distinctive traits.
@@
Our research abides by ethical guidelines for AI research and is committed to privacy, confidentiality, and intellectual property rights. We have ensured that the datasets in our study, which are publicly available, do not house sensitive or personal details. While our study leverages existing resources and generative models, it is important to note that these models can possess biases and may occasionally generate summaries with distortions, biases, or inappropriate content Gooding (2022). We have configured our models to omit potentially harmful or unsafe content to counteract this. While our research aims to enhance meeting summarization to benefit communication and productivity across sectors, we are acutely aware of the ethical challenges posed by AI in this domain. Meeting summarization models must be wielded with respect to privacy and consent, especially when processing sensitive or confidential material. It’s paramount that these models neither violate privacy nor perpetuate harmful biases. As the field evolves, we stress the importance of maintaining these ethical considerations and encourage fellow researchers to uphold them, ensuring that AI advancements in meeting summarization are both beneficial and ethically grounded. An integral aspect of our ethical commitment is reflected in our approach to annotator recruitment and management. The team of annotators, consisting of interns, student assistants, and doctoral students, was meticulously selected through internal channels. This strategy was chosen to uphold a high standard of annotation quality—a quality we found challenging to guarantee through external platforms such as Amazon Mechanical Turk. Ensuring fair compensation, these annotators were reimbursed following institutional guidelines for their respective positions. Further, flexibility in the annotation process was also a priority. Annotators were free to choose their working times and environments to prevent fatigue from affecting their judgment.
Appendix A Annotation process details
Following, we describe the details of our annotation process.
Annotator selection:
Our annotation team comprises four graduate students, officially employed as interns or doctoral candidates through standardized contracts. We select them from a pool of volunteers based on their availability to complete the task without time pressure and their English proficiency (native speakers or C1-C2 certified). This ensures they can comprehend meeting transcripts, human-written gold summaries from QMSum, and model-generated summaries. We aimed for gender balance (two male, two female) and diverse backgrounds, resulting in a team of two computer science students, one psychology student, and one communication science student, aged 24-28.
Preparation:
We prepare a comprehensive handbook for our annotators, detailing the project context and defining challenges and error types. Each definition includes two examples: one with minimal impact (e.g., slight information redundancy) and one with high impact (e.g., repeated information throughout). The handbook explains the two-part rating system: a binary yes/no for the existence of a challenge or error, followed by a 1-5 Likert scale impact/frequency rating if a characteristic is observed as existing. This impact/frequency scoring is only used for the challenges (frequency, 1 low - 5 high) and errors (impact, 1 low - 5 high) produced by the LED model. Annotators are further tasked to provide reasoning for each decision. The handbook does not specify an order for processing errors or challenges. We provide the handbook in English and the annotators’ native languages, using professional translations. The handbook could be used throughout the whole annotation process as a reference.
We set up a five-week timeline for the annotation process, preceded by a one-week onboarding period. The first two weeks feature twice-weekly check-ins with annotators, which are reduced to weekly meetings for the following three weeks. Separate quality checks without the annotators are scheduled weekly. (Note: week refers to a regular working week)
Onboarding:
The onboarding week is dedicated to getting to know the project and familiarizing with the definitions and data. We begin with a kick-off meeting to introduce the project and explain the handbook, mainly focusing on each definition. We note initial questions to revise the handbook potentially. Annotators are provided with 35 samples generated by SLED+BART Ivgi et al. (2022), chosen for their balance of identifiable errors and good-quality summaries while capable of processing the whole meeting. After the first 15 samples, we hold individual meetings to clarify any confusion and update the guidelines accordingly. The remaining 20 samples are then annotated using these updated guidelines. A second group meeting this week addresses any new definitional issues. After the group meeting, we meet individually with annotators to review their work, ensuring quality and understanding of the task and samples. All four annotators demonstrate reliable performance and good comprehension of the task and definitions, judging from the reasoning they provided for each decision and annotation. We computed an inter-annotator agreement score using Krippendorff’s alpha, achieving 0.81, indicating sufficiently high overlap.
Annotation Process:
We continue the annotation process similarly. Each week, we distribute all 35 samples generated by one model to one of the annotators. Consequently, one annotator works through all samples of one model in one week. After five weeks, all samples have been processed by all annotators. Annotators are unaware of the summary-generating model and are given a week to complete their set at their own pace and break times. Quiet working rooms were provided if needed for concentration. To mitigate position bias, the sample order is randomized for each annotator. Annotators can choose their annotation order for each sample and are allowed to revisit previous samples to adapt ratings. To simplify the process, we frame each error type as a question, such as "Does the summary omit crucial information or provide insufficient detail about salient points?" for the missing information error.
Regular meetings are held to address questions on definitions or emerging issues. During the quality checks the authors perform, we look for incomplete annotations, missing explanations, and signs of misunderstanding judging from the provided reasoning. If we find such a lack of quality, the respective annotator will be notified to re-do the annotation. The quality checks are not used to bias annotators in their ratings but to ensure a complete and consistent dataset. After the five weeks, we compute inter-annotator agreement scores (shown in Tables 2 and 3). In case we observe a significant difference across annotators at this point, we have planned a dedicated meeting to discuss such cases with all annotators and a senior annotator to ensure that an understanding issue of the task or definition did not lead to the different ratings.
Annotators spend 43 minutes per sample, completing about seven samples daily.
Handling of scheduling conflicts:
Given that our annotators have other commitments, we anticipate potential scheduling conflicts. We allow flexibility for annotators to complete their samples beyond the week limit if needed, reserving a sixth week as a buffer. Despite these provisions, all annotators completed their assigned samples within the original weekly timeframes.
Appendix B Statistics on Annotation Labels
In Tables 7 and 8, we report statistics of the annotated dataset, showing how many are showing the respective challenge and error type.
Spoken Language Speaker Dynamic Co- reference Discourse Structure Contextual Turn-Taking Implicit Context Data Scarcity Low Inf. Density
# detections 32 (91%) 35 (100%) 35 (100%) 34 (97%) 33 (94%) 5 (14%) 30 (86%) 35 (100%)
Table 7: Statistics on the occurance of different challenges in the QMSum samples used as input. We report total number and corresponding percentage.
MI Red WR IR Hal Inc LI SD LED 31 (89%) 14 (40%) 4 (11%) 7 (20%) 9 (26%) 13 (37%) 3 (9%) 20 (57%) DialogLED 33 (94%) 18 (51%) 4 (11%) 8 (23%) 8 (23%) 18 (51%) 3 (9%) 22 (63%) Pegasus-X 34 (97%) 27 (77%) 14 (40%) 16 (46%) 13 (37%) 23 (66%) 10 (29%) 28 (80%) GPT3.5 34 (97%) 7 (20%) 3 (9%) 1 (3%) 5 (14%) 9 (26%) 9 (26%) 22 (63%) Zephyr 34 (97%) 7 (20%) 3 (9%) 1 (3%) 8 (23%) 11 (31%) 11 (31%) 22 (63%)
Table 8: Statistics on the occurance of different error types in the model-generated summaries. We report total number and corresponding percentage.
Citation for this Paper